In the triangle shown above A = 45°, a = 15 cm, and
c = 8 cm. What is the approximate measure of C?


Answer by jim_thompson5910(35256)

Use the law of sines.

sin(A)/a = sin(C)/c

sin(45)/15 = sin(C)/8

8*sin(45)/15 = sin(C)

8*(sqrt(2)/2)/15 = sin(C)

(4*sqrt(2))/15 = sin(C)

sin(C) = (4*sqrt(2))/15

C = arcsin( (4*sqrt(2))/15 )

C = 22.155626

So the angle C is roughly 22.155626 degrees